# A Survey of Efficient Reasoning for Large Reasoning Models: Language, Multimodality, and Beyond

### v1

@@ 3.1 Length Budgeting
A straightforward method for balancing accuracy and efficiency is explicitly budgeting computational resources during inference. Most LRMs operate within a sequential paradigm, wherein the model reflects on and refines its previously generated thoughts. In this context, managing the length of the reasoning chain serves as an intuitive means of budget allocation. In addition, another effective test-time scaling technique, parallel search Snell et al. (2024), allows for explicit budgeting of candidate numbers to enhance efficiency, which we discuss in Section 3.4.
@@
Previous work Nayab et al. (2024) demonstrates the potential of LLM to adhere to length constraints specified in the prompt. Most existing works utilize this nature to control the generation length directly using specialized prompts. TALE Han et al. (2024) uses zero-shot prompting to estimate an optimal token budget which constrains model generation. Sketch-of-Thought Aytes et al. (2025) enhances LLM reasoning efficiency by reducing verbosity in intermediate reasoning steps with three adaptive paradigms: Conceptual Chaining, Chunked Symbolism, and Expert Lexicon. Apart from imposing an overall budget constraint on the entire reasoning process, recent works have also explored imposing budget restrictions on each reasoning step. Wang et al. (2023b) introduce a hierarchical approach to enhance language model reasoning by incorporating planning tokens at the start of each reasoning step. In contrast to verbose intermediate steps, Xu et al. (2025a) propose Chain-of-Draft, which encourages language models to generate concise, minimal intermediate reasoning steps, rather than token-heavy explanations in traditional CoT.
@@
Instead of integrating length budgeting into prompts, Muennighoff et al. (2025) propose S1, using a budget-forcing strategy where the thinking process is forced to end by appending an end-of-thinking token delimiter, thus directly controlling thinking length. Similarly, Jiang et al. (2025) propose two decoding strategies, including ZeroThink and LessThink, to force the model to start its response without applying any thought or with a short thought process. However, forcibly budgeting length may lead to varying degrees of accuracy degradation Jin et al. (2024); Renze and Guven (2024). To comprehensively investigate the relationship between reasoning length and model performance, Lee et al. (2025) present the first systematic study of diverse compression instructions, revealing a universal trade-off curve between response length and accuracy.
@@
In addition to encouraging concise reasoning, researchers also seek to leverage the query difficulty to dynamically budget generation length. Wang et al. (2024b) propose Difficulty-Adaptive Self-Consistency (DSC), which evaluates the difficulty information of the queries using LLM itself, to dynamically allocate inference resources. Similarly, Dynasor Fu et al. (2024b) allocates compute based on model certainty during multi-path reasoning, assigning more resources to hard queries. On the other hand, Wang et al. (2025e) introduce Thought Switching Penalty (TSP) to discourage premature transitions between thoughts which may cause superficial but lengthy reasoning traces.
@@ 3.2 System Switch
The dual process theory Wason and Evans (1974) explains that human reasoning operates through two systems: an implicit, fast, and intuitive process (System 1) and an explicit, slower, and deliberative process (System 2). Kahneman (2011) extensively explores this framework, highlighting how System 1 enables rapid decision-making but is prone to biases, while System 2 provides logical oversight for complex reasoning and analysis. Building on dual process theory, several studies Su et al. (2024a); Saha et al. (2024); Cheng et al. (2025b); Pan et al. (2024); Sun et al. (2025a) have explored alternating between Systems 1 and 2, carefully allocating computing resources to both systems optimize the balance between reasoning quality and efficiency.
Dualformer Su et al. (2024a) integrates the dual process through a randomized reasoning trace training strategy, which randomly drop certain parts of the reasoning traces. Another approach involves training a switch module to toggle between the two systems Saha et al. (2024); Sun et al. (2025a); Cheng et al. (2025a). These switch systems can be effectively trained using supervised labels in well-defined scenarios such as Maze Navigation. Specifically, Saha et al. (2024) introduce System-1.x, which employs a controller to assess maze difficulty. This allows the model to alternate among different systems based on user-defined parameters for smoother allocation of cognitive resources when addressing sub-goals. Similarly, Sun et al. (2025a) develop a switching adapter that dynamically transitions between Systems 1 and 2 for visual reasoning according to task complexity factors like visual uncertainty and invisibility. Differently, HaluSearch Cheng et al. (2025a) leverages the model performance on specific instances to construct supervised labels, based on which the model learn to switch from System 1 to System 2 in both instance-level and step-level under MCTS (Monte Carlo Tree Search). In addition to training-aware methods, Dyna-Think Pan et al. (2024) uses a training-free dynamic thinking mechanism whereby the model autonomously determines “Slow” reasoning based on generation consistency and complexity of thought processes.
@@ 3.3 Model Switch
Speculative decoding Ryu and Kim (2024) has emerged as a key strategy to accelerate inference by leveraging draft models or early exit mechanisms to propose multiple candidate tokens before verification by the full model. BiLD Kim et al. (2023) utilizes a small, fast model for initial predictions and a larger, more accurate model for corrections, effectively balancing speed and quality through fallback and rollback policies. EAGLE Li et al. (2024e) enhances inference by transitioning speculative sampling from the token level to the feature level. EAGLE-2 Li et al. (2024d) further refines speculative decoding by introducing context-aware dynamic draft trees that adjust token acceptance rates based on confidence scores. In contrast to computing allocation between large and small models, MEDUSA Cai et al. (2024) accelerates large language model inference by incorporating additional decoding heads that predict multiple tokens simultaneously. By integrating a tree-based attention mechanism, it concurrently generates and verifies several candidate continuations, thereby reducing sequential decoding steps. LayerSkip Elhoushi et al. (2024), on the other hand, speeds up inference through layer dropout combined with early exit loss. This allows predictions at shallower layers while also implementing self-speculative decoding for verification purposes.
Another line of work Ong et al. (2025); Wang et al. (2025b); Lu et al. (2023) introduces a routing module to select LLMs for specific prompts given their difficulty and complexity. For example, Lu et al. (2023) propose Zooter, a reward-guided routing method that leverages distilled rewards from training queries to train a specialized routing function. This function accurately directs each query to the LLM with the most pertinent expertise. Ong et al. (2025) introduce RouteLLM, which learns to dynamically route queries between robust and weaker language models, striking an optimal balance between performance and cost effectiveness. Likewise, MixLLM Wang et al. (2025b) enhance query embeddings using tag knowledge, employing lightweight predictors to assess quality and cost per model while leveraging a meta decision maker to select the optimal LLM candidate.
@@ 3.4 Parallel Search
Recent large reasoning models primarily focus on enhancing the efficiency of sequential revision methods, such as o1 or R1 DeepSeek-AI (2025). An additional line of research focuses on enhancing the efficiency of parallel search, another commonly utilized test-time scaling paradigm Snell et al. (2024). Typical methods include majority voting, self-consistency Wang et al. (2023c), and Best-of-N Lightman et al. (2024) which employs a verifier (e.g., voting or a reward model) to select from multiple candidates generated in parallel by the policy model. Expanding the search space, defined as the number of candidates per prompt, consistently improves performance until it plateaus, albeit at an increased computational cost in terms of FLOPs.
To improve the efficiency of parallel search, instead of waiting for all generations to be completed, SBoN Sun et al. (2024a) evaluates partial responses and halts those that are unlikely to yield high-quality completions, achieving comparable performance while substantially reducing computational resource demands. Qiu et al. (2024) propose TreeBoN that combines speculative tree-search with Best-of-N sampling. By generating candidate responses in a hierarchical tree structure, TreeBoN expands high-reward partial responses while pruning low-quality ones early using a weighted implicit reward. A line of work Wang et al. (2025d); Huang et al. (2025a) proposes to replace the external reward model, which is typically the same size as the policy model, further reducing the computational overhead. STBoN Wang et al. (2025d) truncates suboptimal candidates early via identifying the earliest estimation time when samples become distinct, and employing a buffer window along with hidden state consistency. Huang et al. (2025a) distill self-consistency–derived confidence into the model which enables strategies like early stopping and eliminates the need for external reward models. A few works combine sequential revision and parallel search to boost efficiency. Sui et al. (2025) introduce a meta-reasoner with strategies like restarting or refining, using a contextual multi-armed bandit formulation. Li et al. (2025d) propose a recursive approach to revise parallel samples, aligning model performance at test time and achieving comparable results to training-aware methods.
@@ 3.5 Summary and Outlook
Despite the effectiveness of length budgeting methods, research on effectively pruning reasoning traces, specifically precisely eliminating redundant elements, remains limited due to insufficient evaluation of reasoning chain efficacy. Moreover, the integration of inherent model features into length control has been underexplored. For example, a stricter budget may be applied when the model perceives a problem as “trivial”. Similarly, such model-aware adaptiveness can also be applied to speculative decoding, where only tough problems are left for the larger and stronger model. In addition, besides system switch and model switch, model merging Wu et al. (2025a); Lu et al. (2024b) may be a promising direction to balance the task difficulty and reasoning efficiency. Lastly, balancing search depth with search width through parallel search presents a promising approach to significantly reduce inference latency at the expense of increased memory consumption.
@@ 4 Efficient Reasoning with SFT
Supervised fine-tuning (SFT) is a straightforward way to help models learn how to follow the instructions of users Wang et al. (2022b); Zhang et al. (2023b). In this section, we survey existing methods that fine-tune the models to achieve efficient reasoning. As shown in Figure 6, these methods mainly consist of two categories, including training with a compressed reasoning chain and training with tokens in the latent space.
@@ 4.1 Reasoning Chain Compression
To generate a concise reasoning path, Token-Budget-Aware LLM Reasoning Han et al. (2024) first produces the target output by prompting the model with a CoT prompt that includes the optimized token budget, then they train the model with SFT to produce answers that adhere to the token budget. In a different constructing way, Munkhbat et al. (2025) build concise reasoning paths via best-of-N sampling and few-shot conditioning. Then, they apply SFT to distill the length-reduction path into the model. With a more radical approach, Yu et al. (2024a) fine-tune the models to omit the intermediate step of generation for samples that are sufficiently confident.
To eliminate redundant information in the reasoning chain, C3ot Kang et al. (2024) employs GPT-4 Achiam et al. (2023) as a compressor, preserving key information throughout the reasoning process. The model is then fine-tuned to learn the relationship between long and short CoTs. Instead of relying on an external GPT model to filter, LMskip Liu et al. (2024c) focuses on skipping intermediate reasoning steps. To induce step-skipping behavior, a controlled training environment is designed, instructing models to produce reasoning sequences under a step constraint. Subsequently, shorter, yet accurate, reasoning paths are chosen and integrated with complete reasoning paths. This augmented dataset is used to finetune a new model with enhanced step-skipping capabilities. In a different way to select important steps, SPIRIT-FT Cui et al. (2025) identifies key reasoning steps by using perplexity as a metric, a step is deemed critical if its removal significantly increases perplexity. In addition to step skipping, TokenSkip Xia et al. (2025) analyzes token importance in CoT outputs, selectively omitting less important tokens for controllable compression of CoT sequences.
Rather than being limited to a compressed format of reasoning chain, CoT-Valve Ma et al. (2025) fine-tunes a model to generate both long and short reasoning paths. Their approach involves identifying a specific task vector within the parameter space that governs the length of the generated CoT. Significantly, this vector allows for extrapolation, enabling the generation of reasoning chains that are either longer or shorter than those encountered during training.
@@ 4.2 Latent-Space SFT
Another line for efficient reasoning via SFT is latent space reasoning, where explicit CoT steps are gradually replaced by continuous hidden representations. Previous works define latent reasoning as the internal computations in transformers. In this context, intermediate variables in two-hop reasoning can be recovered from hidden states Yang et al. (2024b), while “back-patching” was proposed to intervene in this process Biran et al. (2024). Similarly, Implicit CoT eliminates explicit intermediate steps by directly predicting answers from internal representations rather than generating full token sequences Deng et al. (2024). Moreover, studies on CoT unfaithfulness have revealed that even when a chain-of-thought is generated, the model may internally follow a different latent reasoning process Wang et al. (2023a); Turpin et al. (2023).
@@
Inspired by these insights, Coconut (Chain of Continuous Thought) replaces traditional CoT by using the model’s last hidden state as a continuous representation of reasoning. Instead of generating tokens step-by-step, the hidden state is fed back into the model as input for subsequent reasoning steps. This method is grounded in curriculum learning Wang et al. (2021); Soviany et al. (2022), where the model gradually transitions from generating explicit reasoning steps to operating entirely in the latent space, leading to more efficient reasoning and reducing the token overhead typically associated with CoT. Similarly, CCoT (Compressed Chain of Thought) Cheng and Van Durme (2024) fine-tunes the model to produce compressed representations of reasoning chains instead of full-length sequences. By approximating complete reasoning chains with fewer tokens, CCoT reduces computational cost and enhances throughput, while allowing dynamic adjustment of the performance-efficiency tradeoff during inference.
@@
Building upon this, CODI (Continuous Chain-of-Thought via Self-Distillation) introduces an improvement to the curriculum learning approach by integrating a self-distillation framework. Inspired by previous works on context-based learning and generalized prompt compression Ge et al. (2024); Li et al. (2024f), CODI focuses on aligning the hidden activations of specific tokens between a teacher model using explicit CoT and a student model using implicit CoT Shen et al. (2025c). This alignment spans across all layers of LLM, effectively injecting explicit reasoning into the implicit reasoning process. As a result, CODI enhances performance while addressing the forgetting issue, making it a more robust solution for reasoning tasks.
@@
Token Assorted further improves reasoning efficiency by mixing latent and text tokens Su et al. (2025). By using latent discrete tokens from VQ-VAE, it abstracts the initial reasoning steps, reducing trace length while retaining essential information. This approach results in a 17% reduction in reasoning trace length and enhanced performance on logical and mathematical tasks. SoftCoT takes a different approach to continuous-space reasoning by utilizing an assistant model that generates “soft thought tokens” for the LLM Xu et al. (2025b). These tokens are projected into the LLM’s representation space, enabling efficient reasoning without the need for full model fine-tuning, improving efficiency while preserving pre-trained knowledge.
@@
Furthermore, LightThinker Zhang et al. (2025a) enhances reasoning by dynamically compressing intermediate steps into concise latent representations. This reduces memory usage and computational overhead while maintaining key reasoning information. Similarly, Heima leverages hidden latent representations to reduce verbosity in both text and multimodal tasks Shen et al. (2025a). The Heima Encoder compresses intermediate steps into a single token, and the Heima Decoder reconstructs the reasoning process from these tokens, significantly lowering token usage and improving efficiency. These approaches demonstrate the growing trend of leveraging latent space reasoning to enhance efficiency in LLMs, each offering unique strategies to reduce computational overhead while maintaining or improving reasoning capabilities, paving the way for more scalable and effective models in complex tasks.

### v2

@@ 3.1 Length Budgeting
A straightforward method for balancing accuracy and efficiency is explicitly budgeting computational resources during inference. Most LRMs operate within a sequential paradigm, wherein the model reflects on and refines its previously generated thoughts. In this context, managing the length of the reasoning chain serves as an intuitive means of budget allocation. In addition, another effective test-time scaling technique, parallel search snell2024scaling, allows for explicit budgeting of candidate numbers to enhance efficiency, which we discuss in Section 3.4.
@@
Previous work nayab2024concise demonstrates the potential of LLM to adhere to length constraints specified in the prompt. Most existing works utilize this nature to control the generation length directly using specialized prompts. TALE han2024token uses zero-shot prompting to estimate an optimal token budget which constrains model generation. Sketch-of-Thought aytes2025sketchofthoughtefficientllmreasoning enhances LLM reasoning efficiency by reducing verbosity in intermediate reasoning steps with three adaptive paradigms: Conceptual Chaining, Chunked Symbolism, and Expert Lexicon. Apart from imposing an overall budget constraint on the entire reasoning process, recent works have also explored imposing budget restrictions on each reasoning step. wang2023guiding introduce a hierarchical approach to enhance language model reasoning by incorporating planning tokens at the start of each reasoning step. In contrast to verbose intermediate steps, xu2025chaindraftthinkingfaster propose Chain-of-Draft, which encourages language models to generate concise, minimal intermediate reasoning steps, rather than token-heavy explanations in traditional CoT.
@@
Instead of integrating length budgeting into prompts, muennighoff2025s1 propose S1, using a budget-forcing strategy where the thinking process is forced to end by appending an end-of-thinking token delimiter, thus directly controlling thinking length. Similarly, jiang2025safechain propose two decoding strategies, including ZeroThink and LessThink, to force the model to start its response without applying any thought or with a short thought process. However, forcibly budgeting length may lead to varying degrees of accuracy degradation jin2024impact; renze2024benefits. To comprehensively investigate the relationship between reasoning length and model performance, lee2025llmscompresschainofthoughttoken present the first systematic study of diverse compression instructions, revealing a universal trade-off curve between response length and accuracy.
@@
In addition to encouraging concise reasoning, researchers also seek to leverage the query difficulty to dynamically budget generation length. wang2024make propose Difficulty-Adaptive Self-Consistency (DSC), which evaluates the difficulty information of the queries using LLM itself, to dynamically allocate inference resources. Similarly, Dynasor fu2024efficientlyservingllmreasoning allocates compute based on model certainty during multi-path reasoning, assigning more resources to hard queries. On the other hand, wang2025thoughts introduce Thought Switching Penalty (TSP) to discourage premature transitions between thoughts which may cause superficial but lengthy reasoning traces.
@@ 3.2 System Switch
The dual process theory dualprocess explains that human reasoning operates through two systems: an implicit, fast, and intuitive process (System 1) and an explicit, slower, and deliberative process (System 2). kahneman2011thinking extensively explores this framework, highlighting how System 1 enables rapid decision-making but is prone to biases, while System 2 provides logical oversight for complex reasoning and analysis. Building on dual process theory, several studies su2024dualformercontrollablefastslow; saha2024system1xlearningbalancefast; cheng2025think; pan2024dynathink; sun2025visualagentsfastslow have explored alternating between Systems 1 and 2, carefully allocating computing resources to both systems optimize the balance between reasoning quality and efficiency.
Dualformer su2024dualformercontrollablefastslow integrates the dual process through a randomized reasoning trace training strategy, which randomly drop certain parts of the reasoning traces. Another approach involves training a switch module to toggle between the two systems saha2024system1xlearningbalancefast; sun2025visualagentsfastslow; cheng2025thinkmorehallucinateless. These switch systems can be effectively trained using supervised labels in well-defined scenarios such as Maze Navigation. Specifically, saha2024system1xlearningbalancefast introduce System-1.x, which employs a controller to assess maze difficulty. This allows the model to alternate among different systems based on user-defined parameters for smoother allocation of cognitive resources when addressing sub-goals. Similarly, sun2025visualagentsfastslow develop a switching adapter that dynamically transitions between Systems 1 and 2 for visual reasoning according to task complexity factors like visual uncertainty and invisibility. Differently, HaluSearch cheng2025thinkmorehallucinateless leverages the model performance on specific instances to construct supervised labels, based on which the model learn to switch from System 1 to System 2 in both instance-level and step-level under MCTS (Monte Carlo Tree Search). In addition to training-aware methods, Dyna-Think pan2024dynathink uses a training-free dynamic thinking mechanism whereby the model autonomously determines “Slow” reasoning based on generation consistency and complexity of thought processes.
@@ 3.3 Model Switch
Speculative decoding ryu2024closer has emerged as a key strategy to accelerate inference by leveraging draft models or early exit mechanisms to propose multiple candidate tokens before verification by the full model. BiLD Kim2023SpeculativeDW utilizes a small, fast model for initial predictions and a larger, more accurate model for corrections, effectively balancing speed and quality through fallback and rollback policies. EAGLE li2024eagle enhances inference by transitioning speculative sampling from the token level to the feature level. EAGLE-2 li2024eagle2 further refines speculative decoding by introducing context-aware dynamic draft trees that adjust token acceptance rates based on confidence scores. In contrast to computing allocation between large and small models, MEDUSA Cai2024MedusaSL accelerates large language model inference by incorporating additional decoding heads that predict multiple tokens simultaneously. By integrating a tree-based attention mechanism, it concurrently generates and verifies several candidate continuations, thereby reducing sequential decoding steps. LayerSkip elhoushi-etal-2024-layerskip, on the other hand, speeds up inference through layer dropout combined with early exit loss. This allows predictions at shallower layers while also implementing self-speculative decoding for verification purposes.
Another line of work ong2025routellmlearningroutellms; wang2025mixllmdynamicroutingmixed; lu2023routingexpertefficientrewardguided introduces a routing module to select LLMs for specific prompts given their difficulty and complexity. For example, lu2023routingexpertefficientrewardguided propose Zooter, a reward-guided routing method that leverages distilled rewards from training queries to train a specialized routing function. This function accurately directs each query to the LLM with the most pertinent expertise. ong2025routellmlearningroutellms introduce RouteLLM, which learns to dynamically route queries between robust and weaker language models, striking an optimal balance between performance and cost effectiveness. Likewise, MixLLM wang2025mixllmdynamicroutingmixed enhance query embeddings using tag knowledge, employing lightweight predictors to assess quality and cost per model while leveraging a meta decision maker to select the optimal LLM candidate.
@@ 3.4 Parallel Search
Recent large reasoning models primarily focus on enhancing the efficiency of sequential revision methods, such as o1 or R1 deepseekr1. An additional line of research focuses on enhancing the efficiency of parallel search, another commonly utilized test-time scaling paradigm snell2024scaling. Typical methods include majority voting, self-consistency self_consistency, and Best-of-N verify_step which employs a verifier (e.g., voting or a reward model) to select from multiple candidates generated in parallel by the policy model. Expanding the search space, defined as the number of candidates per prompt, consistently improves performance until it plateaus, albeit at an increased computational cost in terms of FLOPs.
To improve the efficiency of parallel search, instead of waiting for all generations to be completed, SBoN sbon evaluates partial responses and halts those that are unlikely to yield high-quality completions, achieving comparable performance while substantially reducing computational resource demands. treebon propose TreeBoN that combines speculative tree-search with Best-of-N sampling. By generating candidate responses in a hierarchical tree structure, TreeBoN expands high-reward partial responses while pruning low-quality ones early using a weighted implicit reward. A line of work wang2025sampling; huang2025efficienttesttimescalingselfcalibration proposes to replace the external reward model, which is typically the same size as the policy model, further reducing the computational overhead. STBoN wang2025sampling truncates suboptimal candidates early via identifying the earliest estimation time when samples become distinct, and employing a buffer window along with hidden state consistency. huang2025efficienttesttimescalingselfcalibration distill self‐consistency–derived confidence into the model which enables strategies like early stopping and eliminates the need for external reward models. A few works combine sequential revision and parallel search to boost efficiency. sui2025metareasonerdynamicguidanceoptimized introduce a meta-reasoner with strategies like restarting or refining, using a contextual multi-armed bandit formulation. Li2025TestTimePO propose a recursive approach to revise parallel samples, aligning model performance at test time and achieving comparable results to training-aware methods.
@@ 3.5 Summary and Outlook
Despite the effectiveness of length budgeting methods, research on effectively pruning reasoning traces, specifically precisely eliminating redundant elements, remains limited due to insufficient evaluation of reasoning chain efficacy. Moreover, the integration of inherent model features into length control has been underexplored. For example, a stricter budget may be applied when the model perceives a problem as “trivial”. Similarly, such model-aware adaptiveness can also be applied to speculative decoding, where only tough problems are left for the larger and stronger model. In addition, besides system switch and model switch, model merging wu2025unlockingefficientlongtoshortllm; lu2024twin may be a promising direction to balance the task difficulty and reasoning efficiency. Lastly, balancing search depth with search width through parallel search presents a promising approach to significantly reduce inference latency at the expense of increased memory consumption.
@@ 4 Efficient Reasoning with SFT
Supervised fine-tuning (SFT) is a straightforward way to help models learn how to follow the instructions of users wang2022self; zhang2023instruction. In this section, we survey existing methods that fine-tune the models to achieve efficient reasoning. As shown in Figure 6, these methods mainly consist of two categories, including training with a compressed reasoning chain and training with tokens in the latent space.
@@ 4.1 Reasoning Chain Compression
To generate a concise reasoning path, Token-Budget-Aware LLM Reasoning han2024token first produces the target output by prompting the model with a CoT prompt that includes the optimized token budget, then they train the model with SFT to produce answers that adhere to the token budget. In a different constructing way, munkhbat2025selftrainingelicitsconcisereasoning build concise reasoning paths via best-of-N sampling and few-shot conditioning. Then, they apply SFT to distill the length-reduction path into the model. With a more radical approach, yu2024distilling21 fine-tune the models to omit the intermediate step of generation for samples that are sufficiently confident.
To eliminate redundant information in the reasoning chain, C3ot kang2024c3ot employs GPT-4 achiam2023gpt as a compressor, preserving key information throughout the reasoning process. The model is then fine-tuned to learn the relationship between long and short CoTs. Instead of relying on an external GPT model to filter, LMskip liu2024can focuses on skipping intermediate reasoning steps. To induce step-skipping behavior, a controlled training environment is designed, instructing models to produce reasoning sequences under a step constraint. Subsequently, shorter, yet accurate, reasoning paths are chosen and integrated with complete reasoning paths. This augmented dataset is used to finetune a new model with enhanced step-skipping capabilities. In a different way to select important steps, SPIRIT-FT cui2025stepwise identifies key reasoning steps by using perplexity as a metric, a step is deemed critical if its removal significantly increases perplexity. In addition to step skipping, TokenSkip xia2025tokenskip analyzes token importance in CoT outputs, selectively omitting less important tokens for controllable compression of CoT sequences.
Rather than being limited to a compressed format of reasoning chain, CoT-Valve ma2025cot fine-tunes a model to generate both long and short reasoning paths. Their approach involves identifying a specific task vector within the parameter space that governs the length of the generated CoT. Significantly, this vector allows for extrapolation, enabling the generation of reasoning chains that are either longer or shorter than those encountered during training.
@@ 4.2 Latent-Space SFT
Another line for efficient reasoning via SFT is latent space reasoning, where explicit CoT steps are gradually replaced by continuous hidden representations. Previous works define latent reasoning as the internal computations in transformers. In this context, intermediate variables in two-hop reasoning can be recovered from hidden states yang2024largelanguagemodelslatently, while “back-patching” was proposed to intervene in this process biran2024hoppinglateexploringlimitations. Similarly, Implicit CoT eliminates explicit intermediate steps by directly predicting answers from internal representations rather than generating full token sequences deng2024explicit. Moreover, studies on CoT unfaithfulness have revealed that even when a chain-of-thought is generated, the model may internally follow a different latent reasoning process wang2023understandingchainofthoughtpromptingempirical; turpin2023language.
@@
Inspired by these insights, Coconut (Chain of Continuous Thought) replaces traditional CoT by using the model’s last hidden state as a continuous representation of reasoning. Instead of generating tokens step-by-step, the hidden state is fed back into the model as input for subsequent reasoning steps. This method is grounded in curriculum learning wang2021surveycurriculumlearning; soviany2022curriculum, where the model gradually transitions from generating explicit reasoning steps to operating entirely in the latent space, leading to more efficient reasoning and reducing the token overhead typically associated with CoT. Similarly, CCoT (Compressed Chain of Thought) cheng2024compressed fine-tunes the model to produce compressed representations of reasoning chains instead of full-length sequences. By approximating complete reasoning chains with fewer tokens, CCoT reduces computational cost and enhances throughput, while allowing dynamic adjustment of the performance-efficiency tradeoff during inference.
@@
Building upon this, CODI (Continuous Chain-of-Thought via Self-Distillation) introduces an improvement to the curriculum learning approach by integrating a self-distillation framework. Inspired by previous works on context-based learning and generalized prompt compression ge2024incontext; li2024500xcompressorgeneralizedpromptcompression, CODI focuses on aligning the hidden activations of specific tokens between a teacher model using explicit CoT and a student model using implicit CoT shen2025codicompressingchainofthoughtcontinuous. This alignment spans across all layers of LLM, effectively injecting explicit reasoning into the implicit reasoning process. As a result, CODI enhances performance while addressing the forgetting issue, making it a more robust solution for reasoning tasks.
@@
Token Assorted further improves reasoning efficiency by mixing latent and text tokens su2025token. By using latent discrete tokens from VQ-VAE, it abstracts the initial reasoning steps, reducing trace length while retaining essential information. This approach results in a 17% reduction in reasoning trace length and enhanced performance on logical and mathematical tasks. SoftCoT takes a different approach to continuous-space reasoning by utilizing an assistant model that generates “soft thought tokens” for the LLM xu2025softcot. These tokens are projected into the LLM’s representation space, enabling efficient reasoning without the need for full model fine-tuning, improving efficiency while preserving pre-trained knowledge.
@@
Furthermore, LightThinker zhang2025lightthinker enhances reasoning by dynamically compressing intermediate steps into concise latent representations. This reduces memory usage and computational overhead while maintaining key reasoning information. Similarly, Heima leverages hidden latent representations to reduce verbosity in both text and multimodal tasks shen2025efficient. The Heima Encoder compresses intermediate steps into a single token, and the Heima Decoder reconstructs the reasoning process from these tokens, significantly lowering token usage and improving efficiency. These approaches demonstrate the growing trend of leveraging latent space reasoning to enhance efficiency in LLMs, each offering unique strategies to reduce computational overhead while maintaining or improving reasoning capabilities, paving the way for more scalable and effective models in complex tasks.